MC algo 13-13 Calculating Variance

Based on the following information, what is the variance?

State of
Economy
Probability of State
of Economy
Rate of Return if
State Occurs
Recession .24 11.00%
Normal .27 12.50%
Boom .49 23.50%

Multiple Choice

  • .13848

  • .03835

  • .01918

  • .08842

  • .02877

MC algo 13-36 Expected Return

A stock has an expected return of 10.80 percent. Based on the following information, what is the stock's return in a boom state of the economy?

State of Economy Probability of State
of Economy
Rate of Return if
State Occurs
Recession .23 11.1 %
Normal .26 12.6 %
Boom .51 ?

Multiple Choice

  • 20.66%

  • 19.76%

  • 21.56%

  • 17.29%

  • 18.52%
Answer #1

1)

Mean = 0.24*(-0.11) + 0.27*0.125 + 0.49*0.235

Mean = -0.0264 + 0.03375 + 0.11515

Mean = 0.1225 or 12.25%

Variance = [0.24(-0.11 - 0.1225)^2 + 0.27(0.125 - 0.1225)^2 + 0.49(0.235 - 0.1225)^2]

Variance = [0.012974.+ 0.000002 + 0.006202]

Variance = .01918

2)

0.108 = 0.23*(-0.111) + 0.26*0.126 + 0.51*X

0.108 = -0.02553 + 0.03276 + 0.51X

0.10077 = 0.51X

X = 0.1976 or 19.76%